# Ailing Yu1  Lan Yao2  Jingnan Liu1  Zhe Chen1  Jiajun Yin1 Yuan Wang1  Xinhao Liao1  Zhiling Ye1  Ji Li1  Yun Yue1 Hansong Xiao1  Hualei Zhou1  Chunxiao Guo1  Peng Wei1  Jinjie Gu1

### v1

@@
Figure 1: Overall performance of MedReseacher-R1 across three benchmarks. On MedBrowseComp, our MedResearcher-R1-32B achieves state-of-the-art 27.5% accuracy, surpassing o3-deepresearch (25.5%), Gemini-2.5-Pro-deepresearch (25.0%), and significantly outperforming search-only approaches (o3-search: 19.0%, Gemini-2.5-Pro-search: 14.0%). On general deep research tasks, we achieve competitive results on GAIA (53.4 vs. WebSailor-32B’s 53.2) and xBench (54.0 vs. WebSailor-32B’s 53.3).
@@ 1 Introduction
However, medical domain presents unique challenges that current general-purpose deep research agents fail to address adequately. The recently introduced MedBrowseComp benchmark (Chen et al., 2025b) reveals this critical gap: even OpenAI’s o3-deepresearch, the leading proprietary deep research system, achieves only $25.5\%$ accuracy on complex medical queries requiring multi-hop reasoning across medical knowledge sources. We identify two fundamental limitations that contribute to this performance gap: (1) general-purpose agents lack the dense, specialized medical knowledge required for accurate clinical reasoning, and (2) they rely on generic retrieval tools that fail to capture the nuanced relationships in medical information.
@@
First, we develop a novel data synthesis framework that generates training examples of exceptional complexity through a systematic pipeline: We begin by extracting medical entities from over 30 million PubMed abstracts, then apply frequency analysis to identify candidates with occurrence rates below $10^{-6}$ in medical corpora. Through LLM-assisted evaluation, we filter these candidates to select genuinely rare yet clinically significant entities, avoiding both trivial typos and overly common conditions. Around these carefully selected rare medical entities, we construct knowledge graphs to extract longest reasoning chains for multi-hop question generation. This approach creates questions that mirror real medical research challenges and cannot be answered through simple retrieval but require systematic exploration and synthesis across multiple medical information sources.
Figure 2: Comparison of medical reasoning agents.MedReseacher-R1 resolves the Valsartan identification case that defeats general-purpose agents, demonstrating the strength of specialized medical database access and evidence-based reasoning integration.
Second, we introduce proprietary medical domain tools that address retrieval gaps in general systems. As illustrated in Figure 2, while general agents often fail when encountering medical-specific queries, particularly those involving rare diseases or complex chemical compounds, MedReseacher-R1 can iteratively invoke specialized medical tools alongside general-purpose tools to ensure accurate information retrieval. Unlike conventional search engines that rely on general web crawling, our custom-built private medical retrieval engine directly accesses authoritative medical databases including FDA databases, official prescription data, clinical trial registries, and peer-reviewed medical publications. The comparison in Figure 2 demonstrates how MedReseacher-R1 dynamically switch between general and medical-specific tools enables systematic validation of the complete evidence chain: from corporate merger information to heart failure drug development, to chemical composition and mechanism, ultimately ensuring clinical accuracy while avoiding the reasoning errors that plague general-only approaches. The system employs medical ontology-aware ranking to prioritize clinical authority and relevance over general web popularity metrics, effectively combining the breadth of general-purpose search with the precision of domain-specific medical expertise.
@@
Our experimental results validate this approach. The trained model, MedReseacher-R1, achieves $27.5\%$ accuracy on MedBrowseComp, establishing a new state-of-the-art and substantially outperforming both the Qwen2.5-32B baseline and existing deep research systems. Notably, our medical specialization does not compromise general capabilities: on general agent benchmarks (GAIA: 53.4, xBench: 54), MedReseacher-R1maintains competitive performance comparable to GPT-4o-mini. Please refer to Figure 1 for an overview.
@@
2 MedReseacher-R1: Medical Deep Research Agent Framework
@@ 3.1.3 Quality Control and Difficulty Calibration
To ensure that generated questions remain challenging for current systems, we implement adaptive difficulty calibration. Each question is evaluated against GPT-o3 deepresearch and GPT-4, if either model achieves $50\%$ accuracy, the question is automatically regenerated with increased complexity:
@@ 5 Experiments
We evaluate MedReseacher-R1 across both domain-specific and general-purpose benchmarks to assess its effectiveness in complex medical research tasks and its generalization capabilities beyond the medical domain.
@@ 5.1 Benchmarks
  • XBench-DeepSearch (Chen et al., 2025a) is an extensive multi-domain agent evaluation suite that systematically assesses tool-use capabilities across diverse open-domain tasks. The benchmark encompasses a broad spectrum of scenarios including fact-checking, comparative analysis, web browsing-based reasoning, and complex information synthesis tasks, providing a comprehensive evaluation of an agent’s ability to navigate and utilize various tools in real-world problem-solving contexts.
@@ 5.2 Main Results
As shown in Table 1, our tool-augmented agent achieves new state-of-the-art performance on MedBrowseComp benchmark, achieving a pass@1 accuracy of $27.5\%$ and outperforming both previous best agents and the Qwen2.5-32B baseline. The supervised fine-tuning (SFT) stage already delivers notable gains, while subsequent reinforcement learning further improves decision quality and tool orchestration efficiency.
@@ 7 Conclusion
Built on this rich dataset and equipped with a comprehensive training pipeline—including supervised fine-tuning, trajectory masking, and reinforcement learning with specialized medical tools—our agent, MedReseacher-R1 achieves state-of-the-art pass@1 accuracy on MedBrowseComp (27.5%) and demonstrates robust performance on general agent benchmarks. These findings show that MedReseacher-R1 is capable of solving complex medical questions that demand systematic exploration and nuanced evidence synthesis, highlighting its effectiveness as a next-generation deep research agent in the medical domain.

### v2

@@
Figure 1: Overall performance of MedResearcher-R1 across three benchmarks. On MedBrowseComp, our MedResearcher-R1-32B achieves state-of-the-art 27.5% accuracy, surpassing o3-deepresearch (25.5%), Gemini-2.5-Pro-deepresearch (25.0%), and significantly outperforming search-only approaches (o3-search: 19.0%, Gemini-2.5-Pro-search: 14.0%). On general deep research tasks, we achieve competitive results on GAIA (53.4 vs. WebSailor-32B’s 53.2) and xBench (54.0 vs. WebSailor-32B’s 53.3).
@@ 1 Introduction
However, the medical domain presents unique challenges that current general-purpose deep research agents fail to address adequately. The recently introduced MedBrowseComp benchmark (Chen et al., 2025b) reveals this critical gap: even OpenAI’s o3-deepresearch, the leading proprietary deep research system, achieves only $25.5\%$ accuracy on complex medical queries requiring multi-hop reasoning across medical knowledge sources. We identify two fundamental limitations that contribute to this performance gap: (1) general-purpose agents lack the dense, specialized medical knowledge required for accurate clinical reasoning, and (2) they rely on generic retrieval tools that fail to capture the nuanced relationships in medical information.
@@
First, we develop a novel data synthesis framework that generates training examples of exceptional complexity through a systematic pipeline: We begin by extracting medical entities from over 30 million PubMed abstracts, then apply frequency analysis to identify candidates with occurrence rates below $10^{-6}$ in medical corpora. Through LLM-assisted evaluation, we filter these candidates to select genuinely rare yet clinically significant entities, avoiding both trivial typos and overly common conditions. Around these carefully selected rare medical entities, we construct knowledge graphs to extract the longest reasoning chains for multi-hop question generation. This approach creates questions that mirror real medical research challenges and cannot be answered through simple retrieval but require systematic exploration and synthesis across multiple medical information sources.
Figure 2: Comparison of medical reasoning agents. MedResearcher-R1 resolves the Valsartan identification case that defeats general-purpose agents, demonstrating the strength of specialized medical database access and evidence-based reasoning integration.
Second, we introduce proprietary medical domain tools that address retrieval gaps in general systems. As illustrated in Figure 2, while general agents often fail when encountering medical-specific queries, particularly those involving rare diseases or complex chemical compounds, MedResearcher-R1 can iteratively invoke specialized medical tools alongside general-purpose tools to ensure accurate information retrieval. Unlike conventional search engines that rely on general web crawling, our custom-built private medical retrieval engine directly accesses authoritative medical databases, including FDA databases, official prescription data, clinical trial registries, and peer-reviewed medical publications. The comparison in Figure 2 demonstrates how MedResearcher-R1 dynamically switches between general and medical-specific tools, enabling systematic validation of the complete evidence chain: from corporate merger information to heart failure drug development, to chemical composition and mechanism, ultimately ensuring clinical accuracy while avoiding the reasoning errors that plague general-only approaches. The system employs medical ontology-aware ranking to prioritize clinical authority and relevance over general web popularity metrics, effectively combining the breadth of general-purpose search with the precision of domain-specific medical expertise.
@@
Our experimental results validate this approach. The trained model, MedResearcher-R1, achieves $27.5\%$ accuracy on MedBrowseComp, establishing a new state-of-the-art and substantially outperforming both the Qwen2.5-32B baseline and existing deep research systems. Notably, our medical specialization does not compromise general capabilities: on general agent benchmarks (GAIA: 53.4, xBench: 54), MedResearcher-R1maintains competitive performance comparable to GPT-4o-mini. Please refer to Figure 1 for an overview.
@@
2 MedResearcher-R1: Medical Deep Research Agent Framework
@@ 3.1.3 Quality Control and Difficulty Calibration
To ensure that generated questions remain challenging for current systems, we implement adaptive difficulty calibration. Each question is evaluated against OpenAI-o3 deepresearch and GPT-4. If either model achieves $50\%$ accuracy, the question is automatically regenerated with increased complexity:
@@ 5 Experiments
We evaluate MedResearcher-R1 across both domain-specific and general-purpose benchmarks to assess its effectiveness in complex medical research tasks and its generalization capabilities beyond the medical domain.
@@ 5.1 Benchmarks
  • XBench-DeepSearch (Chen et al., 2025a) is an extensive multi-domain agent evaluation suite that systematically assesses tool-use capabilities across diverse open-domain tasks. The benchmark encompasses a broad spectrum of scenarios, including fact-checking, comparative analysis, web browsing-based reasoning, and complex information synthesis tasks, providing a comprehensive evaluation of an agent’s ability to navigate and utilize various tools in real-world problem-solving contexts.
@@ 5.2 Main Results
As shown in Table 1, our tool-augmented agent achieves new state-of-the-art performance on the MedBrowseComp benchmark, achieving a pass@1 accuracy of $27.5\%$ and outperforming both previous best agents and the Qwen2.5-32B baseline. The supervised fine-tuning (SFT) stage already delivers notable gains, while subsequent reinforcement learning further improves decision quality and tool orchestration efficiency.
@@ 7 Conclusion
Built on this rich dataset and equipped with a comprehensive training pipeline—including supervised fine-tuning, trajectory masking, and reinforcement learning with specialized medical tools—our agent, MedResearcher-R1 achieves state-of-the-art pass@1 accuracy on MedBrowseComp (27.5%) and demonstrates robust performance on general agent benchmarks. These findings show that MedResearcher-R1 is capable of solving complex medical questions that demand systematic exploration and nuanced evidence synthesis, highlighting its effectiveness as a next-generation deep research agent in the medical domain.